# Positioning-Aided Channel Estimation for Multi-LEO Satellite Downlink Communications

### v1

@@ IV-C3 Update $\mathbf{W}_{s}$
[…]
@@
[…]
Number of LEO satellites $S$ 8
Number of UTs $U$ 4
Antenna number $N=N_{\text{h}}\times N_{\text{v}}$ $8\times 8$
Antenna radiation gain $G(\cdot)$ $\cos(\cdot)$
V Numerical Results
V-A Simulation Parameters
Figure 4: Illustrative example of the considered scenario.
[…]
@@ V-B3 Downlink Communication
To evaluate the performance of downlink communication, we utilize the CSI obtained through the different channel estimation schemes introduced above to perform distributed beamforming, as summarized in Algorithm 1. The resulting beamformer is tested using ground-truth CSI, yielding the actual sum rate. Additionally, we consider two alternative benchmarks:
  • PAB: This approach is identical to the one introduced for downlink positioning, which relies solely on position information. However, the beamformer at each LEO satellite is applied only to a subset of subcarriers.
  • Single-Satellite Service ($\text{S}^{3}$): In this method, each UT is served by only a single LEO satellite. Specifically, the one providing the strongest signal strength, rather than being served by all satellites within a cluster.
@@ V-C Downlink Positioning
$10^{-1}$$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$er-LEO satellite power budget [dBm]RMSE [m]PAB: LBPAB: MCRBPAB: BiasPAB: CRBVDB: LBVDB: MCRBVDB: BiasVDB: CRB
@@
As shown in Fig. 5, we compare the positioning error bounds and biases under PAB and VDB as functions of the per-satellite power budget. The results reveal that the LB of positioning error without model mismatch (represented by the positioning CRB under perfectly time- and frequency-synchronized satellites) decreases with increasing power. In contrast, the LB with mismatch initially decreases but eventually saturates at a value determined by the bias, which remains independent of the transmit power. This emphasizes the detrimental impact of imperfect time- and frequency-synchronization among LEO satellites on positioning performance [26, 24], potentially leading to overly optimistic performance estimates. Moreover, compared to VDB, PAB achieves a lower LB, with the performance gain being more pronounced at lower (and more practical) transmit power levels. This advantage arises because PAB, by utilizing position information, can allocate more power to the UTs than VDB, underscoring the importance of integrating prior position knowledge into the positioning system.
@@
As illustrated in Fig. 6, we further analyze the impact of varying time and frequency biases among satellite-UT pairs on positioning performance, resulting from imperfect time and frequency synchronization among LEO satellites. Specifically, we present the LB of positioning error (under PAB) as a function of the maximum clock bias and CFO values. The clock bias and CFO for each satellite-UT pair are uniformly distributed between 00 and their respective maximum values. The results reveal that positioning performance deteriorates significantly as the maximum clock bias and CFO increase. A slight clock bias (on the order of milliseconds) or a minor CFO (on the order of $\Delta f/10$) can result in positioning errors of several kilometers, highlighting the critical importance of maintaining precise synchronization among LEO satellites in both time and frequency domains to achieve high-accuracy positioning.
$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{-15}$$10^{-13}$$10^{-11}$$10^{-9}$osition error [m]RMSECRBLBMCRBBias
@@
$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{-13}$$10^{-11}$$10^{-9}$osition error [m]RMSE [s]
@@ V-D Uplink Channel Estimation
In Figs. 7(a) and (b), we evaluate the LB for both channel gain and TOA estimation RMSE using our proposed PACE across varying positioning errors, comparing it with Perfect-PACE (without positioning error-induced model mismatch). The results reveal that when positioning errors are small, PACE’s LB closely matches that of Perfect-PACE for both parameters. However, as positioning errors increase, PACE’s performance deteriorates due to growing model mismatch. For channel gain estimation, we observe that while the MCRB-induced error remains relatively insensitive to positioning errors, the bias component increases proportionally with positioning error, eventually becoming the primary contributor to overall channel gain error. The TOA estimation exhibits different behavior: its bias initially increases with positioning error before reaching a plateau, while the MCRB-induced error continues to grow, ultimately becoming the dominant factor in the total TOA estimation error. These findings underscore the importance of accurate downlink positioning in the first stage of PACE for achieving high-accuracy uplink channel estimation in the second stage.
@@
In Figs. 8(a) and (b), we compare the LB for the RMSE of both channel vector and TOA estimation using the proposed PACE against those obtained under UCE, versus UT’s power budget. Note that PACE estimates both the channel gain $\alpha_{s,u}$ and the TOA $\tau_{s,u}$, whereas UCE estimates the channel vector $\bm{h}_{s,u}=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})\mathbf{a}(\bm{\theta}_{s,u})$ (excluding the TOA-induced phase) in addition to $\tau_{s,u}$. To enable a unified comparison, we transform the LB for channel gain estimation RMSE in PACE into the corresponding LB for channel vector estimation RMSE (using the form $\hat{\bm{h}}_{s,u}=\hat{\alpha}_{s,u}G(\hat{\theta}_{s,u}^{\mathrm{el}})\mathbf{a}(\hat{\bm{\theta}}_{s,u})$) through algebraic manipulation. As shown, the TOA estimation performance of PACE closely aligns with that of UCE, exhibiting only negligible degradation due to positioning error. This alignment occurs because TOA estimation is independent of whether the other unknown parameter is the channel gain or the channel vector. This independence can be readily demonstrated by deriving the CRB for TOA estimation under Perfect-PACE and UCE, which we omit here due to space limitations. In contrast, PACE significantly outperforms UCE in channel vector estimation. This superiority is attributed to PACE’s utilization of position information, which reduces the complexity of estimating a vector (as required by UCE) to that of estimating a scalar, albeit subject to positioning error. Consequently, PACE achieves higher channel estimation accuracy under the same pilot length and power budget.
@@
Figure 9: Sum rate versus per-LEO satellite antenna number.
@@
Figure 10: Sum rate versus per-UT power budget.
V-E Donwlink Communication
In Fig. 9, we compare the downlink communication sum rates under various approaches versus the number of antennas equipped on each LEO satellite. As observed, the sum rate under the proposed PACE significantly outperforms those under UCE and other baselines, reaching a level very close to the sum rate with perfect CSI. This results from PACE’s superior performance in reconstructing the channel accurately. Meanwhile, the sum rate under PAB is even lower than that under $\text{S}^{3}$, underscoring the importance of accurate CSI in delivering the full capability of multi-LEO satellite downlink communication. Furthermore, with an increasing number of antennas on each LEO satellite, the performance gain of the proposed scheme becomes more evident, highlighting its potential in facilitating high-speed LEO satellite communication with massive antenna arrays.
As shown in Fig. 10, we compare the resultant sum rates under various schemes versus the per-UT power budget. It can be seen that the sum rates under PAB and $\text{S}^{3}$ are irrelevant to UTs’ transmit power, while $\text{S}^{3}$ remains superior to PAB throughout the entire range. However, with increasing transmit power from the UT side, the sum rates under PACE and UCE increase and gradually approach those under perfect CSI. This is because the channel estimation accuracy under both schemes increases with higher transmit power, which translates into a sum rate gain. Nevertheless, the sum rate under PACE is significantly always higher than that under UCE (especially notable when the transmit power at the UT is low). This highlights the importance of leveraging position information to obtain accurate CSI when the power budget at the UTs is limited, which is particularly relevant for equipments such as cell phones.
VI Conclusion
This paper presents a novel framework for integrating positioning and communication in multi-LEO satellite systems, leveraging positioning capabilities to enhance channel estimation and thereby facilitate communication. By capitalizing on the distinct timescale variations of position-related parameters and channel gains, we propose a two-timescale frame structure that strategically schedules downlink positioning, uplink channel estimation, and downlink communication. A rigorous MCRB-based analysis characterizes the impact of practical imperfections, including inter-satellite clock bias and CFO, on positioning accuracy. Additionally, we theoretically demonstrate through MCRB analysis how position information improves uplink channel estimation even under inevitable positioning errors. To overcome the limited link budget offered by single-satellite-based downlink communication, we introduce a multi-satellite distributed beamforming strategy that relies on local CSI at each satellite, eliminating the need for centralized coordination. Numerical results validate our theoretical findings. As a direction for future work, developing practical algorithms and validating the framework through experimental testing would be valuable next steps.
Appendix A Derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to $\mathbf{r}_{u}$
@@
$-\frac{\mathbf{v}_{s}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}+\mathbf{v}_{s}^{\mathsf{T}}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\mathbf{I}_{3}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}},$
@@
Appendix B Derivatives of $\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)$ with respect to $\bm{\eta}_{s,u}$

### v4

@@ V-A Simulation Parameters
Central carrier frequency $f_{c}$ $12.7\text{ GHz}$ (Ku band)
Subcarrier spacing $\Delta f$ $120\text{ KHz}$
@@
Power budget at each LEO satellite $50\text{ dBm}$
Power budget at each UT $40\text{ dBm}$
PSD $N_{0}$ $173.855\text{ dBm/Hz}$
Noise figure $F$ $10\text{ dB}$
Number of LEO satellites $S$ 4
Number of UTs $U$ 8
Antenna number $N=N_{\text{h}}\times N_{\text{v}}$ $16\times 16$
Antenna radiation gain $G(\cdot)$ $\sqrt{\frac{3}{4\pi}}\cos(\theta)$[39]
@@ V-B3 Downlink Communication
To evaluate downlink communication performance, we use the estimated CSI from the channel estimation schemes above to perform cooperative beamforming, as outlined in Algorithm 1 (since the satellites only have access to estimated CSI). The resulting beamformer is then tested with ground-truth CSI to obtain the actual sum rate. We also consider two benchmark schemes:
  • PAB: This approach is the same as the one introduced for downlink positioning.
  • No-Cooperation (NC): In this method, each UT is served only by the single LEO satellite with the strongest signal strength, thereby avoiding inter-satellite cooperation. The WMMSE algorithm is applied independently at each satellite to optimize its beamformers based solely on its own CSI for the associated UTs, while signals from other satellites are treated as interference.
@@ V-C Downlink Positioning
[…]
@@
As shown in Fig. 3, we compare the positioning error bounds and biases under PAB and VDB with different per-satellite power budget. The results reveal that the LB of positioning error without model mismatch, represented by the positioning CRB under perfectly time- and frequency-synchronized satellites, decreases as the transmit power increases. In contrast, when synchronization mismatch is present, the LB initially decreases but eventually saturates due to the persistent positioning bias, which remains unaffected by the transmit power. This highlights the detrimental impact of imperfect time- and frequency-synchronization among LEO satellites on positioning performance [37, 35], potentially leading to overly optimistic performance evaluations. Furthermore, PAB consistently outperforms VDB by achieving a lower LB, with the performance gain being particularly significant (over an order of magnitude) at lower yet more practical transmit power levels. This advantage stems from PAB’s ability to leverage position information to allocate more power to the UTs than VDB, underscoring the importance of incorporating prior position knowledge into the design of positioning systems.
@@
In Fig. 4, we analyze the impact of varying clock bias and CFO among satellite–UT pairs on the LB of positioning error (under PAB), where both impairments are assumed to be uniformly distributed between 0 and their respective maximum values. The results show that positioning performance degrades significantly as either factor increases. In particular, even a clock bias on the order of tens to hundreds of nanoseconds or a CFO of several kHz can lead to positioning errors exceeding $100\text{ m}$. These findings highlight the comparable severity of both impairments and underscore the critical importance of mitigating mismatches in both the time and frequency domains to ensure high-accuracy LEO positioning.
$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{-6}$$10^{-4}$$^{-2}1001$$^{0}$Position error [m]Normalized RMSEPerfect-PACE: CRBPACE: LBPACE: MCRBPACE: Bias
@@
$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$^{5}1010$$^{-10}$$10^{-8}$$10^{-6}$Position error [m]Normalized RMSE
@@ V-D Uplink Channel Estimation
In Figs. 5(a) and (b), we evaluate the normalized RMSE, obtained by dividing the RMSE by the ground-truth parameters, for both channel gain and TOA estimation using our proposed PACE, under varying positioning errors. The results are compared with those of Perfect-PACE, which assumes no model mismatch due to positioning errors. The results reveal that when positioning errors are small, PACE’s LB closely matches that of Perfect-PACE for both parameters. However, as positioning errors increase, PACE’s performance deteriorates due to growing model mismatch. For channel gain estimation, we observe that while the MCRB-induced error remains relatively insensitive to positioning errors, the bias component increases proportionally with positioning error, eventually becoming the primary contributor to the overall channel gain error. The TOA estimation exhibits different behavior: its bias initially increases with positioning error before reaching a plateau, whereas the MCRB-induced error continues to grow, ultimately becoming the dominant factor in the total TOA estimation error. Nevertheless, the results indicate that channel estimation degrades noticeably only when the positioning error becomes impractically large (e.g., exceeding $1\text{ km}$). This observation highlights the robustness of PACE to positioning imperfections in LOS-dominant scenarios with relatively high satellite altitudes.
@@
In Figs. 6(a) and (b), we compare the LB for the RMSE of both channel vector and TOA estimation using the proposed PACE against those obtained under UCE, versus UT’s power budget. Note that PACE estimates both the channel gain $\alpha_{s,u}$ and the TOA $\tau_{s,u}$, whereas UCE estimates the channel vector $\bm{h}_{s,u}=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})\mathbf{a}(\bm{\theta}_{s,u})$ (excluding the TOA-induced phase) in addition to $\tau_{s,u}$. To enable a unified comparison, we transform the LB for channel gain estimation RMSE in PACE into the corresponding LB for channel vector estimation RMSE (using the form $\hat{\bm{h}}_{s,u}=\hat{\alpha}_{s,u}G(\hat{\theta}_{s,u}^{\mathrm{el}})\mathbf{a}(\hat{\bm{\theta}}_{s,u})$) through algebraic manipulation. As shown, the TOA estimation performance of PACE closely aligns with that of UCE, given that negligible positioning error is incurred in the positioning stage. This alignment occurs because TOA estimation is independent of whether the other unknown parameter is the channel gain or the channel vector. This independence can be readily demonstrated by deriving the CRB for TOA estimation under Perfect-PACE and UCE, which we omit here due to space limitations. In contrast, PACE significantly outperforms UCE in channel vector estimation. This superiority is attributed to PACE’s utilization of position information, which reduces the complexity of estimating a vector (as required by UCE) to that of estimating a scalar, albeit subject to positioning error. Consequently, PACE achieves higher channel estimation accuracy under the same pilot length and power budget.
V-E Donwlink Communication
In Fig. 7, we compare the downlink communication sum rates achieved by various approaches as a function of the number of antennas deployed on each LEO satellite. It is observed that Algorithm 1 combined with PACE significantly outperforms UCE and other baseline methods, achieving sum rates that closely approach those obtained with perfect CSI. This superior performance is attributed to PACE’s enhanced channel reconstruction accuracy and the benefits of multi-LEO satellite cooperation. In contrast, the sum rate under NC, whether using PACE or UCE, remains consistently higher than that achieved by PAB, even though PAB leverages inter-satellite cooperation. This highlights the critical role of accurate CSI in enhancing downlink communication efficiency, beyond what can be achieved using position information alone. Moreover, as the number of antennas per LEO satellite increases, the performance advantage of the proposed scheme, i.e., the combination of PACE and Algorithm 1, becomes increasingly pronounced, demonstrating its potential for enabling high-throughput LEO satellite communications with large-scale antenna arrays.
4416166464256256055101015152020Antenna number at each LEO satelliteSum rate [Mbps]PACE + Algo. 1UCE + Algo. 1Perfect CSI + Algo. 1PABPACE + NCUCE + NC
Figure 7: Sum rate versus per-LEO satellite antenna number.
202030304040505060607070055101015152020Per-UT power budget [dBm]Sum rate [Mbps]PACE + Algo. 1UCE + Algo. 1Perfect CSI + Algo. 1PABPACE + NCUCE + NC
Figure 8: Sum rate versus per-UT power budget.
As shown in Fig. 8, we compare the sum rates under various schemes versus the per-UT power budget. It can be observed that the sum rate under PAB is independent of the UTs’ transmit power. In contrast, as the UTs’ transmit power increases, the sum rates achieved by Algorithm 1 and NC also increase. This is because higher transmit power improves channel estimation accuracy, which in turn reduces the mismatch between the channel parameters used to optimize beamforming and the ground-truth ones, thereby enhancing the resultant sum rate. Specifically, the sum rates achieved by Algorithm 1 significantly outperform those under NC and PAB, and gradually approach the performance achieved with perfect CSI. This demonstrates the superiority of multi-LEO satellite cooperative beamforming. Moreover, the use of PACE results in a substantially higher sum rate compared to UCE, especially when the UTs’ transmit power is low, due to its superior channel estimation accuracy. This highlights the importance of leveraging position information to obtain accurate CSI when the power budget at the UTs is limited, an aspect particularly relevant for power-constrained devices such as cell phones.
VI Conclusion
This paper presents a novel framework for integrating positioning and communication in multi-LEO satellite systems, leveraging positioning capabilities to enhance channel estimation and thereby facilitate communication. By capitalizing on the distinct timescale variations of position-related parameters and channel gains, we propose a two-timescale frame structure that strategically schedules downlink positioning, uplink channel estimation, and downlink communication. A rigorous MCRB-based analysis characterizes the impact of practical imperfections, including inter-satellite clock bias and CFO, on positioning accuracy. Additionally, we theoretically demonstrate through MCRB analysis how position information improves uplink channel estimation even under inevitable positioning errors. To overcome the limited link budget offered by single-satellite-based downlink communication, we introduce a multi-satellite cooperative beamforming optimization scheme to serve each UT simultaneously with multiple satellites within a cluster while reducing complexity. Numerical results validate our theoretical findings.
Beyond the positioning-aided communication emphasized in this work, studying the reverse direction is equally important for future research. For example, inter-UT sidelink communication and ranging enable information exchange among proximal UTs and provide independent timing and distance references. Such mechanisms are promising for mitigating inter-satellite clock biases and thereby improving overall positioning accuracy.
Appendix A MCRB Basics
This subsection briefly recaps the fundamentals of MCRB. For a parameter $\bar{\bm{\theta}}\in\mathbb{R}^{M\times 1}$, the LB matrix of the mean squared error of a mismatched estimator is provided by[38]
@@ Appendix A MCRB Basics
and
@@
Appendix B Derivatives of $\tilde{\mathbf{f}}\left(\mathbf{r}_{u}\right)$ with respect to $\mathbf{r}_{u}$
@@
$-\frac{\mathbf{v}_{s}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)^{\mathsf{T}}+\mathbf{v}_{s}^{\mathsf{T}}\left(\mathbf{p}_{u}-\mathbf{p}_{s}\right)\mathbf{I}_{3}}{\lambda\|\mathbf{p}_{u}-\mathbf{p}_{s}\|^{3}}$
@@
Appendix C Derivatives of $\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)$ with respect to $\bm{\eta}_{s,u}$
@@
Appendix D Low-Complexity Method for Solving (IV-C3)
D-A Extracting Per-Satellite Subproblem
To reduce complexity of solving (IV-C3), we instead adopt a BCD strategy, optimizing beamformers for each satellite sequentially while fixing the others. For the $s$-th satellite, the subproblem depends only on $\mathbf{w}_{s}=[\mathbf{w}_{s,1}^{\mathsf{T}},\ldots,\mathbf{w}_{s,U}^{\mathsf{T}}]^{\mathsf{T}}$ and can be reformulated as
$\mathop{\min}\limits_{\mathbf{w}_{s}}\;\;\;$ $-2\sum_{u=1}^{U}\Re\left\{\omega_{u}\mu_{u}\mathbf{h}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}\right\}$
$+\sum_{u=1}^{U}\sum_{m=1}^{U}\omega_{u}\left|\mu_{u}\right|^{2}\left|\mathbf{h}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}+\Omega_{s,u,m}\right|^{2}$
${\rm{s.t.}}\;\;\;$ $\left\|\mathbf{w}_{s}\right\|^{2}\leq\frac{P_{s}}{K},$ (47)
where $\Omega_{s,u,m}=\sum_{s^{{}^{\prime}}\neq s}^{S}\mathbf{h}_{s^{{}^{\prime}},u}^{\mathsf{T}}\mathbf{w}_{s^{{}^{\prime}},m}$.
Let $\mathbf{E}_{u}=[\mathbf{0}_{N\times N},\ldots,\underbrace{\mathbf{I}_{N}}_{\text{The }u\text{-th matrix}},\ldots,\mathbf{0}_{N\times N}]\in\mathbb{R}^{N\times NU}$ be the selection matrix. Through algebraic manipulation, the above can be recast as a convex QCQP
$\mathop{\min}\limits_{\mathbf{w}_{s}}\;\;\;$ $\mathbf{w}_{s}^{\mathsf{H}}\mathbf{Q}_{s}\mathbf{w}_{s}-2\Re\left\{\mathbf{b}_{s}^{\mathsf{H}}\mathbf{w}_{s}\right\}$
${\rm{s.t.}}\;\;\;$ $\left\|\mathbf{w}_{s}\right\|^{2}\leq\frac{P_{s}}{K},$ (48)
where
$\mathbf{R}_{s}=\sum_{u=1}^{U}\omega_{u}\left|\mu_{u}\right|^{2}\mathbf{h}_{s,u}^{*}\mathbf{h}_{s,u}^{\mathsf{T}}\succeq\mathbf{0}_{N\times N},$ (49a)
$\mathbf{Q}_{s}=\mathbf{I}_{U}\otimes\mathbf{R}_{s},$ (49b)
$\mathbf{b}_{s}=\sum_{u=1}^{U}\omega_{u}\mu_{u}^{*}\mathbf{E}_{u}^{\mathsf{T}}\mathbf{h}_{s,u}^{*}-\sum_{u=1}^{U}\sum_{m=1}^{U}\omega_{u}\left|\mu_{u}\right|^{2}\Omega_{s,u,m}\mathbf{E}_{m}^{\mathsf{T}}\mathbf{h}_{s,u}^{*}.$
D-B Exploiting Strong Duality
It is straightforward to observe that the above problem is a convex QCQP involving only $\mathbf{w}_{s}$. Rather than relying on a convex solver such as CVX, which uses interior-point methods and incurs a complexity of $\mathcal{O}((NU)^{3})$, we exploit the problem’s strong duality to further reduce computational complexity. Specifically, the Lagrangian of the problem is
$\mathcal{L}\left(\mathbf{w}_{s},\lambda\right)=\mathbf{w}_{s}^{\mathsf{H}}\mathbf{Q}_{s}\mathbf{w}_{s}-2\Re\left\{\mathbf{b}_{s}^{\mathsf{H}}\mathbf{w}_{s}\right\}-\lambda\left(\left\|\mathbf{w}_{s}\right\|^{2}-\frac{P_{s}}{K}\right),$
where $\lambda\geq 0$ is the Lagrangian multiplier associated with the power constraint. Then, the optimal solutions of (D-A) can be derived through examining its Karush–Kuhn–Tucker (KKT) conditions as follows
$\left(\mathbf{Q}_{s}+\lambda\mathbf{I}_{NU}\right)\mathbf{w}_{s}$ $=\mathbf{b}_{s},$ (50a)
$\left\|\mathbf{w}_{s}\right\|$ $=\sqrt{\frac{P_{s}}{K}},$ (50b)
$\lambda$ $\geq 0.$ (50c)
It follows from (50a) that
$\mathbf{w}_{s}=\left(\mathbf{Q}_{s}+\lambda\mathbf{I}_{NU}\right)^{-1}\mathbf{b}_{s}.$ (51)
Therefore, we must choose $\lambda$ such that $\left\|(\mathbf{Q}_{s}+\lambda\mathbf{I}_{NU})^{-1}\mathbf{b}_{s}\right\|=\sqrt{P_{s}/K}$. Define
$g\left(\lambda\right)=\left\|(\mathbf{Q}_{s}+\lambda\mathbf{I}_{NU})^{-1}\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}.$ (52)
[…]
Figure 9: Run time and sum rate comparison between the QCQP-based and proposed low-complexity algorithms: (a) Run time versus LEO satellite number; (b) Sum rate versus LEO satellite number; (c) Run time versus antenna number; (d) Sum rate versus antenna number; (e) Run time versus UT number; (f) Sum rate versus UT number.
D-C Reducing Matrix Inversion Complexity
We note that (52) involves a $NU$-dimensional matrix inversion, which still incurs a computational complexity of $\mathcal{O}((NU)^{3})$ per iteration of the line search. To reduce this complexity, we exploit the structure $\mathbf{Q}_{s}=\mathbf{I}_{U}\otimes\mathbf{R}_{s}$. Let $\mathbf{b}_{s}=[\mathbf{b}_{s,1}^{\mathsf{T}},\ldots,\mathbf{b}_{s,U}^{\mathsf{T}}]^{\mathsf{T}}$, where $\mathbf{b}_{s,u}\in\mathbb{C}^{N}$. We then have
$g\left(\lambda\right)$ $=\left\|\left(\mathbf{I}_{U}\otimes\mathbf{R}_{s}+\lambda\mathbf{I}_{NU}\right)^{-1}\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}$
$=\left\|\left(\mathbf{I}_{U}\otimes\left(\mathbf{R}_{s}+\lambda\mathbf{I}_{N}\right)\right)^{-1}\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}$
$=\left\|\left(\mathbf{I}_{U}\otimes\left(\mathbf{R}_{s}+\lambda\mathbf{I}_{N}\right)^{-1}\right)\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}$
$=\left\|\left(\mathbf{I}_{U}\otimes\left(\mathbf{U}\left(\mathbf{\Lambda}+\lambda\mathbf{I}_{N}\right)^{-1}\mathbf{U}^{\mathsf{H}}\right)\right)\mathbf{b}_{s}\right\|^{2}-\frac{P_{s}}{K}$
$=\sum_{u=1}^{U}\left\|\mathbf{U}\left(\mathbf{\Lambda}+\lambda\mathbf{I}_{N}\right)^{-1}\mathbf{U}^{\mathsf{H}}\mathbf{b}_{s,u}\right\|^{2}-\frac{P_{s}}{K}$
$=\sum_{u=1}^{U}\sum_{n=1}^{N}\frac{\left|\varpi_{s,u,n}\right|^{2}}{\left(\lambda_{n}+\lambda\right)^{2}}-\frac{P_{s}}{K},$ (53)
where $\bm{\varpi}_{s,u}=[\varpi_{s,u,1},\ldots,\varpi_{s,u,N}]^{\mathsf{T}}=\mathbf{U}^{\mathsf{H}}\mathbf{b}_{s,u}$. Here, $\mathbf{R}_{s}$ is decomposed as $\mathbf{R}_{s}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\mathsf{H}}$ via eigenvalue decomposition, where $\mathbf{U}\in\mathbb{C}^{N\times N}$ is a unitary matrix and $\mathbf{\Lambda}=\text{diag}[\lambda_{1},\ldots,\lambda_{N}]$ is the diagonal eigenvalue matrix. Then, (51) is reformulated as
$\mathbf{w}_{s}=\left(\mathbf{I}_{U}\otimes\left(\mathbf{U}\left(\mathbf{\Lambda}+\lambda\mathbf{I}_{N}\right)^{-1}\mathbf{U}^{\mathsf{H}}\right)\right)\mathbf{b}_{s},$ (54)
which eliminates the need for an $NU$-dimensional matrix inversion and thereby reduces the computational complexity of updating beamformers.
D-D Analyzing Complexity
The above reformulation shows that eigenvalue decomposition of the matrix $\mathbf{R}_{s}$ needs to be performed only once with complexity $\mathcal{O}(N^{3})$, rather than repeatedly inverting an $NU$-dimensional matrix for each $\lambda$. The Golden-section search method, which has a linear convergence rate, finds an $\epsilon$-solution within $\mathcal{O}(\log(1/\epsilon))$ iterations. Since each iteration only requires evaluating a one-dimensional function, the associated complexity can generally be neglected. Consequently, the eigenvalue decomposition step dominates the overall complexity of solving (D-A). As (D-A) is solved sequentially for $s=1,\ldots,S$, the total complexity of solving (IV-C3) is reduced to $\mathcal{O}(SN^{3})$, compared with the original $\mathcal{O}((SNU)^{3})$.
To further demonstrate the superiority of the proposed algorithm, Fig. 9 compares the run time and sum rate of the QCQP-based and proposed low-complexity algorithms under varying numbers of LEO satellites, antennas per LEO, and UTs. The results show that the proposed algorithm, which decomposes the beamformer update subproblem via BCD and exploits duality and line search, achieves orders-of-magnitude complexity reduction while maintaining identical sum-rate performance compared with the original QCQP-based update. This is notable since the QCQP formulation yields the optimal subproblem solution, whereas the proposed BCD-based structure converges only to a stationary point. These results validate the efficiency of the proposed low-complexity design.